the length of a hypotenuse is 20 cm. The length of one of the legs is 15 cm. What is the length of the other leg?
Answer by Boreal(15235)

15^2+x^2=20^2=400
225+x^2=400
x^2=175
x=sqrt (175)=13.23 cm.